OpenStudy (anonymous):

How many solutions does the following system of equations have? 2y=5x+4 y=3x+2 a) infinitely many b) zero c) one d) two

OpenStudy (zzr0ck3r):

this is two equations of two different lines IF they are the same line THEN they will have INFINITE solutions IF they are NOT the same line but have the SAME slope, then there are NO solutions IF they have different slopes then there is ONE solution so lets get them to look more like the equations of a line we are used to y = mx+b one is already in that form y = 3x+2 the other one 2y=5x+4 divide every term by 2 y = (5/2) x + 4 so we have two different lines with NON equal slopes what is the answer?

OpenStudy (anonymous):

Then, there would be one solution?

OpenStudy (kc_kennylau):

yes

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

2 will never be the answer....

OpenStudy (kc_kennylau):

@zzr0ck3r \(x^2+3x+2=0\) how many solutions?

OpenStudy (zzr0ck3r):

9-4*2>0 so 2

OpenStudy (zzr0ck3r):

this is not two equations of lines though

OpenStudy (kc_kennylau):

i love this xD http://screencast.com/t/jz087EjjYt

OpenStudy (zzr0ck3r):

two lines can never cross each other exactly twice in Euclid space

OpenStudy (kc_kennylau):

a curve is a line

OpenStudy (zzr0ck3r):

semantics, linear or affine

OpenStudy (kc_kennylau):

never mind i'll stop making fun of you xD

OpenStudy (zzr0ck3r):

when someone says an equation of a line, they are talking about an affine function:)

OpenStudy (zzr0ck3r):

also y = mx+b is only linear if b = 0 so y = 3x+3 is not a linear equation despite what everyone thinks

OpenStudy (zzr0ck3r):

only y = mx is linear

OpenStudy (zzr0ck3r):

in other words....every linear equation goes through the origin in R^2